OpenStudy (anonymous):

After 2 years, Deion earned $270 in simple interest from a CD into which he initially deposited $6000. What was the annual interest rate of the CD? A. 2.25% B. 3.25% C. 9.75% D. 6.5%

OpenStudy (anonymous):

@Orion1213

OpenStudy (perl):

a CD is a certificate of deposit, its a type of loan to a bank thats supposed to generate more interest than just money in a bank account

OpenStudy (perl):

than just money sitting in a bank account*

OpenStudy (perl):

in any case, we can use the formula I = P * R * T

OpenStudy (perl):

270 = 6000*R*2

OpenStudy (perl):

R = 270/(12000) = .0225 R = 2.25%